Hi, I'm hoping someone can please help me! I'm having issues trying to figure out this problem. Your help is greatly appreciated! Thank you!!!
.
Base of polyhedral Vertices Faces Edges
Triangle 4vertices 4faces 6edges
Square ____vertices ____faces 8edges
Pentagon ____vertices 6faces ____edges
Hexagon 7vertices ____faces ____edges
What type of polyhedral is represented in the table?
If n represents the number of sides of the polyhedra base, then write an equation for the number of vertices, faces, and edges of the polyhedral.

Answer by Jk22(389)

Supposing the problem were to build a polyhedron with the same regular polygon as faces :


Taking Euler's formula for polyhedron, we have :

V - E + F = 2 (vertex minus edge + face number equals 2)

suppose the faces were the same base polygon with n side (n vertices)

then :

n*F/2 = E (each side of the F polygons (with n sides) are shared by 2 faces)

n*F/3 = V (each vertex of the n*F is shared by 3 faces)


the equation becomes : n*F/3 - n*F/2 + F = 2

or (6-n)F = 12

F = 12/(6-n), E = 6n/(6-n), V = 4n/(6-n)

triangle : n = 3 : V = 12/3 = 4, E = 18/3 = 6, F = 12/3 = 4

square : n = 4 : V = 16/2 = 8, E = 24/2 = 12, F = 12/2 = 6

pentagon : n = 5 : V = 20, E = 30, F = 12

hexagon : n = 6, all are infinite, this covers the plane.

if the vertex is shared by m faces : V = n*F/m

n*F/m - n*F/2 + F = 2

(2m + 2n - nm)F = 4m

F = 4m/(m(2-n)+2n)

if n=3 (base triangle), F = 4m/(6-m),

m-----3----4-----5
4m----12---16----20
6-m---3----2-----1
-----------------
F-----4----8-----20
tetrahedron, octahedron and icosahedron,

if n=4 (base square), F = 4m/(8-2m)

m=3, F = 12/2 = 6 (cube) (only 3 faces per vertex possible)


n=5 (base pentagon), F = 4m/(10-3m), => m=3 dodecahedron



other solutions : the number of faces per vertex are mixed (like complementary fullerenes) : suppose base as a triangle

suppose n1 and n2 are the possible number of faces per vertex, with 3F/k vertices n2-connected, and 3F(1-1/k) n1-connected

F(n2(6-n1) + 6/k*(n1-n2)) = 4n1n2

for n1=5, n2=3, : F=12, k=6 : 10 vertices 5-connected, 2 vertices 3-connected

for n1=3, n2=5, F=20, k=2, 10 vertices 3-connected, and 10 5-connected




if the polyhedron is made out of 2 faces' type, there are also other cases :
polyhedra called fullerenes
(hexagons and other polygons, or other mix of polygons).
the formula becomes : n1*(6-s1) + n2*(6-s2) = 12, where ni are the numbers of regular polygons with si sides.

This gives solution like :
heptagons (2,3,4,6) and pentagons (14, 15, 16, 18)

4 heptagons and 8 squares

or heptagons (6) and triangles (6) should be able to build a polyhedron ?